Stan sold his snow plow and accessories for $352.00. If he got seven times more for the snow plow, then the accessories, how much did he make for the snow plow?
Please show work.

Answer by shree840(260)

Let x be price of accessories
snow plough is 7x
x+7x=352
8x=352
x=352/8=44
snow plough =7*44=308